OpenStudy (anonymous):

How to factorise x^2+2xy+y^2 ?? Thanks :)

OpenStudy (anonymous):

do you know how to factor:\[a^2+2ab+b^2\]

OpenStudy (anonymous):

your equation is of that form

OpenStudy (anonymous):

you can write:\[a^2+2ab+b^2\]as:\[(a+b)(a+b)\]

OpenStudy (anonymous):

can you apply that to your equation?

OpenStudy (anonymous):

@Brianna98 ?

OpenStudy (anonymous):

well in your case its:\[x^2+2xy+y^2\]

OpenStudy (anonymous):

(the a and b is replaced by x an y)

OpenStudy (anonymous):

so you can factor it as: \[(x+y)(x+y)\]

OpenStudy (anonymous):

did u even understand what i said?

OpenStudy (anonymous):

@Brianna98

OpenStudy (unklerhaukus):

and obscure method would be to split the middle term up x^2+2xy+y^2 =x^2+xy + xy+y^2 then factor two terms at a time =x(x+y) + y(x+y) =

OpenStudy (unklerhaukus):

an*

OpenStudy (anonymous):

thats great too @UnkleRhaukus

OpenStudy (anonymous):

@mihirb Thank you so much!! Makes a lot more sense now :)

OpenStudy (anonymous):

And thank you @UnkleRhaukus :)